# GENIE: Watermarking Graph Neural Networks for Link Prediction

### v1

@@ Abstract
Graph Neural Networks (GNNs) have advanced the field of machine learning by utilizing graph-structured data, which is ubiquitous in the real world. GNNs have applications in various fields, ranging from social network analysis to drug discovery. GNN training is strenuous, requiring significant computational resources and human expertise. It makes a trained GNN an indispensable Intellectual Property (IP) for its owner. Recent studies have shown GNNs to be vulnerable to model-stealing attacks, which raises concerns over IP rights protection. Watermarking has been shown to be effective at protecting the IP of a GNN model. Existing efforts to develop a watermarking scheme for GNNs have only focused on the node classification and the graph classification tasks.
To the best of our knowledge, we introduce the first-ever watermarking scheme for GNNs tailored to the Link Prediction (LP) task. We call our proposed watermarking scheme GENIE (watermarking Graph nEural Networks for lInk prEdiction). We design GENIE using a novel backdoor attack to create a trigger set for two key methods of LP: (1) node representation-based and (2) subgraph-based. In GENIE, the watermark is embedded into the GNN model by training it on both the trigger set and a modified training set, resulting in a watermarked GNN model. To assess a suspect model, we verify the watermark against the trigger set. We extensively evaluate GENIE across 3 model architectures (i.e., SEAL, GCN, and GraphSAGE) and 7 real-world datasets. Furthermore, we validate the robustness of GENIE against 11 state-of-the-art watermark removal techniques and 3 model extraction attacks. We also demonstrate that GENIE is robust against ownership piracy attack. Our ownership demonstration scheme statistically guarantees both False Positive Rate (FPR) and False Negative Rate (FNR) to be less than $10^{-6}$.
@@ 1 Introduction
Deep learning has revolutionized a vast number of Machine Learning (ML) tasks, such as image classification [1], video processing [2], and natural language processing-related tasks [3]. The data utilized in these tasks is typically represented in the Euclidean space. Recently, an increasing number of applications have started utilizing data that is represented in non-Euclidean spaces, such as graphs. Some of the notable applications include social network analysis [4], traffic prediction [5], anomaly detection [6], and drug discovery [7]. The innate complexity of the graphs has posed significant challenges in adapting deep learning techniques to these applications. Fortunately, recent advancements have led to the development of Graph Neural Networks (GNNs) [8], which help in adapting graph data to various machine-learning tasks. GNNs have bridged the gap between deep learning and graphs data structures, enabling the application of sophisticated machine learning models to a wide range of graph-based problems.
GNNs demonstrate remarkable performance across three fundamental tasks: node classification [9] for predicting node properties, graph classification [10] for assigning labels to graphs, and Link Prediction (LP) [11] for inferring the existence of connections between nodes. Developing state-of-the-art GNN models is a formidable task that requires substantial computational resources, domain expertise, and intellectual efforts. It often requires large-scale datasets for training, extensive hyperparameter tuning, and meticulous fine-tuning of model architectures and parameters. The inherent complexity of graph-structured data and the need to effectively capture and propagate structural information pose significant challenges during the training process. GNNs are prone to model stealing attacks [12, 13, 14] that enable an adversary to reproduce the core functionality of the model, even without the knowledge of the victim model’s architecture or the training data distribution. It leads to violations of Intellectual Property (IP) rights. Thus, it is essential to establish ownership and protect the IP rights of GNN model’s owner. Watermarking [15, 16, 17] is a promising technique to verify the ownership of Deep Neural Network (DNN) models. It safeguards the substantial resources invested in developing the model, thereby fostering a safer environment for sharing and collaborating on such models.
Motivation: The area of GNN watermarking remains largely unexplored. Existing works [18, 19] have primarily focused on watermarking GNNs for node classification and graph classification tasks, leaving an opportunity to explore watermarking techniques for LP task. The LP task forms the foundation for many essential operations for Alibaba [20], Facebook [21], and Twitter [22]. Given the significance of LP tasks in real-world scenarios, it is important to secure the ownership and IP rights of the underlying GNN models. Failing to do so could lead to potential misuse of these models by malicious actors as well as financial losses and competitive disadvantages. There are different approaches to perform the complex task of LP, e.g., heuristic methods [23], embedding-based techniques [24], subgraph-based methods [25], and node representation-based models [8]. Creating a watermarking strategy that can handle all these different approaches while still maintaining the model’s performance could be challenging. In this work, we present, to the best of our knowledge, the first-ever watermarking scheme for LP tasks on GNNs. We call our scheme GENIE (watermarking Graph nEural Networks for lInk prEdiction).
@@
    We propose GENIE, a novel approach to watermark GNNs for LP task with minimal utility loss. In particular, GENIE is designed for two key methods of LP in GNNs (viz., node representation-based and subgraph-based methods).
@@
    We propose a novel ownership demonstration scheme having a competitive statistical guarantee of the False Positive Rate (FPR) and False Negative Rate (FNR) being less than $10^{-6}$.
@@
    We perform an extensive evaluation of GENIE using 3 model architectures on 7 real-world datasets. Moreover, we empirically assess the robustness of GENIE against 11 state-of-the-art watermark removal techniques and 3 model extraction attacks.
@@
    We also show that GENIE is computationally efficient and robust to ownership piracy tests.
Organization: The remainder of this paper is organized as follows. §2 presents a summary of the related works and the relevant background knowledge. We elucidate the threat model and our system’s architecture in §3 and §4, respectively. §5 presents our results. §6 highlights the potential limitations of our proposed solution and suggests possible future research directions. Finally, §7 concludes the paper.
@@ 2.1 Graph Neural Networks
Formally, a graph $\mathcal{G}$ is defined as a two-tuple $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges of the graph. We can describe $\mathcal{G}$ using a binary adjacency matrix $\mathbf{A}$ of dimension $|\mathcal{V}|\times|\mathcal{V}|:\mathbf{A}[u,v]=1,\forall(e_{uv})\in\mathcal{E}$ and 00 otherwise. Here, $e_{uv}$ denotes an edge of $\mathcal{G}$ between nodes $u$ and $v$. Modern GNNs take a graph represented by its adjacency matrix $\mathbf{A}$ along with the node feature matrix defined as $\mathbf{X}=[\mathbf{x}_{1};~{}\mathbf{x}_{2};~{}\ldots~{};~{}\mathbf{x}_{|\mathcal{V}|}]_{|\mathcal{V}|\times d}$, where $\mathbf{x}_{i}=[x_{i1}~{}x_{i2}~{}\ldots~{}x_{id}]_{1\times d}$ is the row vector for the $i^{th}$ node for some arbitrary ordering of $\mathcal{V}$ and $d$ is the number of features present in all the nodes as input. Depending on the downstream task, GNN will also take a label vector $\mathbf{y}$ containing the labels. The labels may be of: (1) nodes, in case of node classification; (2) graphs, in case of graph classification; or (3) edges, in case of LP. LP is a semi-supervised/unsupervised task, where an incomplete set of edges $\mathcal{E}_{train}\subset\mathcal{E}$ is given while the goal is to infer the existence of missing edges \$\mathcal{E}\backslash\mathcal{E}_{train}$. Since only positive links [^1] We consider a formulation of LP task, where positive links signify the existing links in a graph, and our goal is to infer the set of missing links at test time. are given, one needs to sample negative links from \$\mathcal{E}\backslash\mathcal{E}_{train}$ making the task unsupervised. The number of samples is generally chosen to be equal to $|\mathcal{E}_{train}|$. As our work focuses on LP, we assume GNNs to get the label vector $\mathbf{y}$ containing labels for both positive links and negative links. GNNs update the representation of each node by iteratively aggregating representations of its neighbors. After $k$ iterations of aggregation, each node’s representation captures structure and feature information within its $k$-hop neighborhood. Typically, the update step for node representations $\textbf{H}^{(k)}$ at $k^{th}$ layer can be represented as:
@@
where $\textbf{W}^{(k)}$ is the learnable weight matrix of the GNN at the $k^{th}$ layer. While training the GNN model, a softmax layer can be used that takes node representations obtained by the GNN for downstream tasks, such as graph classification, LP, etc. The primary difference between different GNN architectures lies in the implementation of the AGGREGATE function. We employ GCN [8], GraphSAGE [26], and SEAL [25] in our work. Each of these is briefly explained below.
SEAL is a widely used LP framework that uses subgraph features. SEAL converts each link into a graph by extracting a local subgraph around it, and then learns a function that maps each subgraph to its corresponding link’s existence using a GNN. In practice, SEAL takes a LP task and treats it as a binary classification problem for graphs. Following the work in SEAL [25], we also use DGCNN [27] model in our work.
GCN uses spectral methods to apply convolutions on graphs. Its $k^{th}$ layer can be summarised as:
$\mathbf{h}_{v}^{(k)}=f^{(k)}\left(\mathbf{W}^{(k)}\cdot\sum_{\begin{subarray}{c}u\in\mathcal{N}(v)\end{subarray}}\frac{\mathbf{h}_{u}^{(k-1)}}{\left|\mathcal{N}(v)\right|}\right),$ (2)
where $f^{(k)}$ is activation function at the $k^{th}$ layer, $\mathbf{h}_{v}^{(k)}$ is the node representation of $v\in\mathcal{V}$ at the $k^{th}$ layer, and $\mathcal{N}(v)$ is the set of nodes directly connected to ${v}$.
GraphSAGE uses spatial-based neighborhood aggregation and sampling to generate node embeddings. Its $k^{th}$ layer can be summarised as:
[…]
where $\text{AGG}_{k}$ is the AGGREGATE function used in the $k^{th}$ layer and CONCAT is the concatenation operation. GraphSAGE [26] proposes three different aggregator functions, i.e., MEAN, MAX-pool, and LSTM. We use the MEAN aggregator for GraphSAGE in our work.
2.2 Link prediction
LP is the task of identifying whether a connection between two nodes is likely to exist. LP studies can broadly be organized into two categories:
Non-DNN based approaches: These include heuristic-based approaches (e.g., CN [23], RA [28]) that predict the presence of a link based on a score, as well as embedding-based approaches (e.g., Matrix Factorization [29], Node2Vec [24]) that learn embeddings via neighborhood information in unsupervised settings to predict the existence of a link.
DNN based approaches: These include using GNN (e.g., GCN [8], GraphSAGE [26] to learn node embeddings via multi-hop neighborhood information, as well as using GNNs with subgraph-features (e.g., SEAL [25]). We focus on DNN based approaches for LP tasks. In particular, we work with two widely adopted GNN architectures (i.e., GCN and GraphSAGE), as well as a subgraph-based framework (i.e., SEAL) for LP.
@@ 2.3 Backdoor attacks and watermarking
DNN watermarking is a method to detect a stolen or extracted model, called the surrogate model, from the original model [35]. DNN watermarking methods can broadly be classified into two categories: (1) static watermarking, which involves embedding a secret signature or mark (called watermark) directly into the model weights during the training phase, and (2) dynamic watermarking, which involves embedding the watermark by changing the weights during the training phase such that the behavior of the model is affected [15]. Static watermarking requires white-box access for the purpose of ownership demonstration. Thus, it is often considered to be unrealistic when the stolen model is available only in a black-box way (e.g., via a prediction API). GENIE classifies as a dynamic watermarking method as it uses backdoored inputs as watermark inputs. Watermarking of GNN serves as proof of ownership and can aid in identifying the model’s ownership.
@@ 2.4 Related works
Our proposed work explores watermarking GNNs in a black-box setting, where the watermark verification process involves using only the target model’s prediction labels (i.e., soft) and not its weights. Adi et al. [16] introduce backdoor-based watermarking for Neural Networks (NNs). Subsequently, numerous watermarking schemes [36, 34, 37, 38] for DNNs have been developed. However, only a few works [18, 19] focus on GNNs owing to its unique structural properties. Zhao et al. [18] introduce watermarking to GNNs using random graphs, but their approach is limited to node classification task. Xu et al. [19] extend the work [18] to node and graph classification tasks using backdoor attacks.
LP, a task with an increasingly large number of applications across diverse domains [11, 39, 40, 41], requires robust mechanisms to protect IP rights and ownership of the underlying models. While there are existing works on backdoor attacks for LP [42, 43], they are not suitable for watermarking purposes. Specifically, the approach proposed by Zheng et al. [42] involves training a separate surrogate model on the same dataset to create the trigger set, which makes it computationally inefficient for large graphs. On the other hand, Dai et al. [43] make the assumption that node features of the graph are binary, which significantly limits the scope of their work. Unlike the works [42, 43], GENIE does not require training a separate surrogate model or node features in binary range. To the best of our knowledge, our work is the first to propose a watermarking scheme for GNNs for LP task. TABLE I presents the description of symbols used in our work.
@@ 3 Threat model
We consider a setting where a model owner $\mathcal{O}$ has trained a GNN model $\mathcal{M}_{own}$ on training data for LP task, and $\mathcal{O}$ deploys it as a service. An adversary $\mathcal{A}$ obtains [^2] The exact mechanism by which $\mathcal{A}$ obtains $\mathcal{M}_{own}$ is out of the scope of our work. a copy of $\mathcal{M}_{own}$, modifies it to create a new model $\mathcal{M}_{adv}$, and $\mathcal{A}$ sets up a competing service based on $\mathcal{M}_{adv}$. Our goal is to protect IP rights over $\mathcal{M}_{own}$ and help confirm whether $\mathcal{M}_{adv}$ has plagiarized from $\mathcal{M}_{own}$. As the model owner, we have full access to the architecture, training data, and the training process. We evaluate our scheme against model extraction attacks (cf. §5). Next, we discuss the philosophy of our approach, watermarking requirements, and the assumptions of our system.
Philosophy of our approach: Our goal is to be able to confirm whether $\mathcal{A}$ has plagiarized from $\mathcal{M}_{own}$ or not with minimal FNR and FPR. Instead of deploying $\mathcal{M}_{own}$ directly as a service, we first watermark it (and obtain $\mathcal{M}_{wm}$) so that even if $\mathcal{A}$ tries to make changes to a copy of $\mathcal{M}_{wm}$, we should still be able to identify the ownership of the plagiarized model (i.e, $\mathcal{M}_{adv}$). To embed the watermark, we train the model on training data $\mathcal{D}_{train}$ and a custom (i.e., to which only $\mathcal{O}$ has access) dataset $\mathcal{D}_{wm}$. We confirm the ownership of a model based on the performance on $\mathcal{D}_{wm}$. If $\mathcal{M}_{adv}$ performs well on $\mathcal{D}_{wm}$, we say $\mathcal{M}_{adv}$ is plagiarized; otherwise, it is not. We must construct $\mathcal{D}_{wm}$ in such a way that $\mathcal{M}_{own}$ can learn it, and the watermark must survive against various attempts by $\mathcal{A}$ to remove the watermark.
Watermarking requirements: We now define the typical requirements [16, 15] for a watermarking scheme.
@@
    Capacity: It refers to the amount of information a watermark can carry. There are two common approaches, viz., zero-bit and multi-bit watermarking [17]. GENIE is primarily designed for ownership verification, i.e., to check if the watermark is present or not. Thus, it does not carry any extra information, making it a zero-bit watermarking approach.
  7. 7.
@@
Assumption: We work under the assumption of $\mathcal{O}$ (plaintiff; who is the true owner of a model) accusing $\mathcal{A}$ (defendant; who has stolen the model from $\mathcal{O}$). In general, the plaintiff and the defendant can both be guilty or innocent. Therefore, ensuring low FPR and FNR [^3] The term false alarm and missed detection probabilities may also be used here as in the work [15]. is a must. In the rest of the paper, our focus will be on the setting, where $\mathcal{A}$ (as a defendant) is guilty and $\mathcal{O}$ (as a plaintiff) is innocent. There are mainly two assumptions we make about $\mathcal{A}$: (1) $\mathcal{A}$ is limited in its knowledge and computational capacity, otherwise it makes stealing the model non-lucrative for $\mathcal{A}$; and (2) $\mathcal{A}$ will make $\mathcal{M}_{adv}$ available via a prediction API that outputs soft labels as a publicly available ML service, potentially disrupting $\mathcal{O}$’s business edge as a niche [^4] Our work does not consider $\mathcal{A}$ that uses $\mathcal{M}_{adv}$ for private use in the same spirit as in media watermarking schemes, which require access to the stolen media as a pre-requisite for ownership demonstration [44].. To evaluate GENIE under various robustness tests, we make a stronger assumption regarding the availability of data, i.e., $\mathcal{A}$ has access to publicly available test data. For demonstration of ownership, the two main assumptions are: (1) a judge $\mathcal{J}$ that (a) ensures confidentiality and correctness of the data, code, etc. submitted to it, and (b) truthfully verifies the output of $\mathcal{M}_{adv}$ evaluated against $\mathcal{D}_{wm}$; and (2) $\mathcal{O}$ and $\mathcal{A}$ will abide by the laws when a dispute is raised. $\mathcal{J}$ plays an integral role in delivering the final verdict of ownership. In practice, $\mathcal{J}$ is typically implemented via a trusted execution environment [45, 46].

### v3

@@ Abstract
Graph Neural Networks (GNNs) have become invaluable intellectual property in graph-based machine learning. However, their vulnerability to model stealing attacks when deployed within Machine Learning as a Service (MLaaS) necessitates robust Ownership Demonstration (OD) techniques. Watermarking is a promising OD framework for Deep Neural Networks, but existing methods fail to generalize to GNNs due to the non-Euclidean nature of graph data. Previous works on GNN watermarking have primarily focused on node and graph classification, overlooking Link Prediction (LP).
In this paper, we propose Genie (watermarking Graph nEural Networks for lInk prEdiction), the first-ever scheme to watermark GNNs for LP. Genie creates a novel backdoor for both node-representation and subgraph-based LP methods, utilizing a unique trigger set and a secret watermark vector. Our OD scheme is equipped with Dynamic Watermark Thresholding (DWT), ensuring high verification probability ($>99.99\%$) while addressing practical issues in existing watermarking schemes. We extensively evaluate Genie across 4 model architectures (i.e., SEAL, GCN, GraphSAGE and NeoGNN) and 7 real-world datasets. Furthermore, we validate the robustness of Genie against 11 state-of-the-art watermark removal techniques and 3 model extraction attacks. We also show Genie’s resilience against ownership piracy attacks. Finally, we discuss a defense strategy to counter adaptive attacks against Genie.
@@ 1 Introduction
Graph Neural Networks (GNNs) have bridged the gap between deep learning and graph data structures, enabling the application of sophisticated machine learning models to a wide range of graph-based problems. Developing State-of-the-Art (SotA) GNN models is a formidable task that requires substantial computational resources, domain expertise, and intellectual efforts. It makes a trained model an indispensable Intellectual Property (IP). However, when deployed as a Machine Learning as a Service (MLaaS), these well-trained models become accessible to both legitimate users and adversaries. Since GNNs are prone to model stealing attacks [1, 2, 3], an adversary can reproduce the core functionality of the model; even without the knowledge of the victim model’s architecture or the training data distribution. It leads to violations of IP rights, making Ownership Demonstration (OD) and protecting IP rights of the model owner essential.
Watermarking [4, 5, 6] is a promising OD framework of Deep Neural Network (DNN) models. There are mainly two settings considered for OD in DNN watermarking [4]: (1) white-box setting, where DNN model parameters have to be known; and (2) black-box setting, where only the final output of the DNN model is accessible. With the rise of MLaaS platforms (where model parameters are unknown), white-box watermarking schemes are considered to be unrealistic (cf. §2.3). Therefore, most watermarking schemes are implemented via backdoors for black-box verification [7, 5]. Existing OD methods employed in watermarking schemes are fraught with problems: (1) lack of efficiency, where OD entails training upto 400 independent models (e.g., in work [7]); (2) lack of theoretical or statistical assurance (e.g., in works [8, 7]); and (3) lack of generalizability to other watermarking schemes (e.g., in work [9]). Most watermarking schemes are also known to fail against adaptive attacks [10]. These problems in watermarking schemes greatly limit their real-world applicability.
Furthermore, DNN watermarking schemes do not generalize to GNN models due to the non-Euclidean nature of graph-structured data. This fundamental incompatibility has left GNN watermarking as a largely unexplored field. The existing works on GNN watermarking [11, 12, 13] have primarily focused on only two of the three tasks of GNNs: node classification and graph classification. These methods fail to trivially generalize to the important task of Link Prediction (LP) as: 1) it differs fundamentally from node/graph classification due to its focus on pairs of nodes or subgraphs rather than individual nodes or graphs; and 2) designing watermarks for LP requires addressing diverse LP methods (node-representation, subgraph-based), which are uncommon for node and graph classification tasks. Consequently, the domain of LP watermarking has remained unexplored. Though there are works on LP backdoor [14, 15], we find them to be impractical for watermarking (cf. §2.4). LP forms the foundation for key operations for Alibaba [16], Facebook [17], and Twitter [18]. Given the significance of LP and GNNs in the real-world, it is important to secure the ownership and IP rights of model owners.
There are mainly two GNN-based approaches to perform LP, viz., subgraph-based methods [19], node representation-based methods [20]. Creating a watermarking scheme that can handle all these different approaches while maintaining the model’s performance is challenging. In this work we present, to the best of our knowledge, the first-ever scheme to watermark GNNs for LP. We call our scheme Genie (watermarking Graph nEural Networks for lInk prEdiction).
Genie creates a novel backdoor for both node-representation and subgraph-based LP methods to watermark GNNs. In particular, Genie operates by first creating a unique trigger set, i.e., the watermark [^1] The term trigger set is used in the context of backdoor attacks, while the term watermark or watermarking dataset ($\mathcal{D}_{wm}$) is used in the context of DNN watermarking. In our work, we use these two terms interchangeably. for both the LP methods. The trigger set is created by random sampling of objects (nodes, in case of node-representation based methods, and links, in case of subgraph-based methods) followed by their corruption with a secret watermark vector (w). A novel watermark embedding method is then used to inject the watermark into the GNN model. We compare our watermark embedding method with other SotA methods in §5.3. Finally, Genie addresses the problem of real-world applicability present in existing watermarking schemes by: (1) using a novel watermark threshold mechanism, called Dynamic Watermark Thresholding (DWT); and (2) dealing with potential adaptive attacks.
@@
    We propose Genie, the first-ever scheme to watermark GNNs for LP with minimal utility loss. In particular, Genie creates a novel backdoor for two key methods of LP in GNNs (viz., node representation-based and subgraph-based methods) (cf. §4).
@@
    We propose a novel watermark embedding method to inject the watermark into the GNN model (cf.  §4.2).
@@
    We propose an OD scheme equipped with a novel watermark threshold mechanism called DWT, which makes the OD procedure efficient, general and with provable statistical confidence assures the False Positive Rate (FPR) and False Negative Rate (FNR) to be less than $10^{-6}$ (cf. §4.3.2).
@@
    We perform an extensive evaluation of Genie using 4 model architectures on 7 real-world datasets (cf. §5). Moreover, we empirically assess the robustness of Genie against 11 SotA watermark removal techniques and 3 model extraction attacks (cf. §5.4).
  5. 5.
    We also show that Genie is robust to ownership piracy tests (cf. §5.5) and computationally efficient (cf. §5.7).
  6. 6.
    We further consider adaptive attackers (who are aware of the working of Genie) and present a defense to mitigate this potential vulnerability (cf. §5.6).
Organization: The remainder of this paper is organized as follows. §2 presents a summary of the related works and the relevant background knowledge. We elucidate the threat model and our system’s architecture in §3 and §4, respectively. §5 presents our evaluation results. Finally, §6 concludes the paper.
@@ 2.1 Graph Neural Networks
Formally, a graph $\mathcal{G}$ is defined as a two-tuple $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges of the graph. We can describe $\mathcal{G}$ using a binary adjacency matrix $\mathbf{A}$ of dimension $|\mathcal{V}|\times|\mathcal{V}|:\mathbf{A}[u,v]=1,\forall(e_{uv})\in\mathcal{E}$ and 00 otherwise. Here, $e_{uv}$ denotes an edge of $\mathcal{G}$ between nodes $u$ and $v$. Modern GNNs take a graph represented by its adjacency matrix $\mathbf{A}$ along with the node feature matrix defined as $\mathbf{X}=[\mathbf{x}_{1};~{}\mathbf{x}_{2};~{}\ldots~{};~{}\mathbf{x}_{|\mathcal{V}|}]_{|\mathcal{V}|\times d}$, where $\mathbf{x}_{i}=[x_{i1}~{}x_{i2}~{}\ldots~{}x_{id}]_{1\times d}$ is the row vector for the $i^{th}$ node for some arbitrary ordering of $\mathcal{V}$ and $d$ is the number of features present in all the nodes as input. GNNs update the representation of each node by iteratively aggregating representations of its neighbors. After $k$ iterations of aggregation, each node’s representation captures structure and feature information within its $k$-hop neighborhood. Typically, the update step for node representations $\textbf{H}^{(k)}$ at $k^{th}$ layer can be represented as:
@@
where $\textbf{W}^{(k)}$ is the learnable weight matrix of GNN at the $k^{th}$ layer. The primary difference between different GNN architectures lies in the implementation of AGGREGATE function. While training the GNN model, a softmax layer can be used that takes node representations obtained by the GNN for downstream tasks, e.g., graph classification, LP.
2.2 Link Prediction (LP)
LP is the task of identifying whether a connection between two nodes is likely to exist. Formally, given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, let $\mathcal{U}$ represent the set of all node pairs possible on $\mathcal{V}$. LP aims to identify potential or missing links in $\mathcal{E}_{u}=\mathcal{U}-\mathcal{E}$, where $\mathcal{E}_{u}$ consists of node pairs not currently connected in $\mathcal{G}$. Our focus is on GNN-based LP methods, which are broadly organized based on the input structure into two categories [21]:
Node-representation based: Such methods (e.g., GCN [20]) take $(\mathbf{A},~{}\mathbf{X})$ of $\mathcal{G}$ as input, obtain node representations, and make predictions.
Subgraph based: Such methods (e.g., SEAL [19]) take as input $(\mathbf{A}_{k},~{}\mathbf{X}_{k})$ of the $k$-hop subgraph $\mathcal{G}_{k}$ surrounding each link, and converts LP into a binary graph classification task.
@@ 2.3 Backdoor attacks and watermarking
DNN watermarking is a method to detect a stolen or extracted model, called the surrogate model, from the original model [10]. DNN watermarking methods can broadly be classified into two categories: (1) static watermarking, which involves embedding a secret signature or mark (called watermark) directly into the model weights during the training phase; and (2) dynamic watermarking, which involves embedding the watermark by changing the weights during the training phase such that the behavior of the model is affected [4]. Static watermarking is often considered to be impractical [7] as: (1) it requires white-box access for the purpose of OD; and (2) it fails when extracted models have different architecture or hyperparameters. Genie classifies as a dynamic watermarking method as it uses backdoored inputs as watermark inputs.
@@ 2.4 Related works
Adi et al. [5] introduce backdoor-based watermarking for Neural Networks (NNs). Subsequently, numerous watermarking schemes [7, 26, 9, 27] for DNNs have been developed. However, only a few works [11, 12, 13] focus on GNNs owing to its unique structural properties. Zhao et al. [11] introduce watermarking to GNNs using random graphs, but their approach is limited to node classification task. Xu et al. [12] extended the work [11] to node and graph classification tasks using backdoor attacks. However, none of the previous works focus on watermarking GNNs for LP.
LP, a task with an increasingly large number of applications across diverse domains [28, 29, 30, 31], requires robust mechanisms to protect IP rights and ownership of the underlying models. While there are existing works on backdoor attacks for LP [14, 32, 15], they are not suitable for watermarking purposes. Specifically, the approach proposed by Zheng et al. [14] involves training a separate surrogate model on the same dataset to create the trigger set, which makes it computationally inefficient for large graphs. Chen et al. [32] proposed an LP backdoor attack targeting discrete-time dynamic graphs, which is inapplicable to generic graphs (the focus of our work) due to different datasets and models. On the other hand, Dai et al. [15] make the assumption that node features of the graph are binary, which significantly limits the scope of their work. Unlike the works [14, 15], Genie does not require training a separate surrogate model or node features in binary range. TABLE I presents the description of symbols used in our work.
@@ 3 Threat model
We consider a setting where a model owner $\mathcal{O}$ has trained her GNN model $\mathcal{M}_{own}$ on training data for LP and deploys it as an MLaaS. An adversary $\mathcal{A}$ obtains [^2] The exact mechanism by which $\mathcal{A}$ obtains $\mathcal{M}_{own}$ is out of the scope of our work. a copy of $\mathcal{M}_{own}$, modifies it to create a new model $\mathcal{M}_{adv}$, and sets up a competing MLaaS based on $\mathcal{M}_{adv}$. Our goal is to protect IP rights over $\mathcal{M}_{own}$ and help confirm whether $\mathcal{M}_{adv}$ is plagiarized from $\mathcal{M}_{own}$. As the model owner, we have full access to the architecture, training data, and the training process. Next, we discuss the philosophy of Genie, watermarking requirements, and the assumptions of our system.
Philosophy of Genie: Our goal is to be able to confirm whether $\mathcal{A}$ has plagiarized from $\mathcal{M}_{own}$ or not with minimal FNR and FPR. Instead of deploying $\mathcal{M}_{own}$ directly as an MLaaS, we first watermark it (and obtain $\mathcal{M}_{wm}$) so that even if $\mathcal{A}$ tries to make changes to a copy of $\mathcal{M}_{wm}$, we should still be able to identify the ownership of the plagiarized model (i.e, $\mathcal{M}_{adv}$). To embed the watermark, we train the model on training data $\mathcal{D}_{train}$ and a secret, custom dataset $\mathcal{D}_{wm}$. We confirm the ownership of $\mathcal{M}_{adv}$ based on its performance on $\mathcal{D}_{wm}$. If $\mathcal{M}_{adv}$ performs well on $\mathcal{D}_{wm}$, we say $\mathcal{M}_{adv}$ is plagiarized; otherwise, it is not. We must construct $\mathcal{D}_{wm}$ in such a way that a GNN model can learn the watermark, and survive various attempts by $\mathcal{A}$ to remove it.
Watermarking requirements: We now define the typical requirements [4] for a watermarking scheme.
@@
Assumption: We work under the assumption of $\mathcal{O}$ (plaintiff; who is the true owner of a model) accusing $\mathcal{A}$ (defendant; who has stolen the model from $\mathcal{O}$). In general, both the plaintiff and the defendant can either be malicious or innocent. Therefore, ensuring low FPR and FNR is a must. In the rest of the paper, our focus will be on the setting, where $\mathcal{A}$ (as a defendant) is malicious and $\mathcal{O}$ (as a plaintiff) is innocent, since we employ judge-generated $\mathcal{D}_{wm}$ (cf. §4.4) to deal with malicious plaintiff and innocent defendant. There are mainly two assumptions we make about $\mathcal{A}$: (1) $\mathcal{A}$ is limited in its knowledge and computational capacity, otherwise it makes stealing the model non-lucrative for $\mathcal{A}$; and (2) $\mathcal{A}$ will make $\mathcal{M}_{adv}$ available via a prediction API that outputs soft labels as an MLaaS, potentially disrupting $\mathcal{O}$’s business edge as a niche [^3] Our work does not consider $\mathcal{A}$ that uses $\mathcal{M}_{adv}$ for private use in the same spirit as in media watermarking schemes, which requires access to the stolen media as a pre-requisite for OD [33].. To evaluate Genie under various robustness tests, we make a stronger assumption regarding the availability of data, i.e., $\mathcal{A}$ has access to publicly available test data. For demonstration of ownership, the two main assumptions are: (1) a judge $\mathcal{J}$ exists that (a) ensures confidentiality and correctness of the data, code, etc. submitted to it, (b) calculates the watermark threshold, and (c) truthfully verifies the output of $\mathcal{M}_{adv}$ evaluated against $\mathcal{D}_{wm}$; and (2) $\mathcal{O}$ and $\mathcal{A}$ will abide by the laws when a dispute is raised. $\mathcal{J}$ plays an integral role in delivering the final verdict of ownership. In practice, $\mathcal{J}$ is typically implemented via a trusted execution environment or a trusted compute base [34, 35].